OpenStudy (anonymous):

How do you simplify [1+(r/n]^1/4

OpenStudy (brooke2353):

break it down

OpenStudy (mimi_x3):

\[ \left ( 1+ \frac{r}{n} \right )^{\frac{1}{4}}\]

OpenStudy (anonymous):

@Mimi_x3 that's not simplified: it's how I wrote it

OpenStudy (mimi_x3):

\[x^{\frac{1}{4}}\] is another way of saying \[\sqrt[4]{x}\]

OpenStudy (mimi_x3):

so \[\left ( 1+ \frac{r}{n} \right )^{\frac{1}{4}}=\sqrt[4]{1+ \frac{r}{n}}\]

OpenStudy (anonymous):

@Mimi_x3 thank you:)

OpenStudy (mimi_x3):

np :)

OpenStudy (mimi_x3):

@dan815 pls medal + fan me

OpenStudy (dan815):

i dont fan just anyone

OpenStudy (mimi_x3):

Wow how do I win u over?

OpenStudy (dan815):

send batsheva my love

OpenStudy (dan815):

by feb14th

OpenStudy (dan815):

I want her to be my valentine

OpenStudy (mimi_x3):

lmao was just talking to her :P Why u sooo creepy like that

OpenStudy (anonymous):

@dan815 so what IS your story?